Circular No. 8008

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002hl IN NGC 3665 T. Boles, Coddenham, England, reports the discovery of an apparent

supernova on an unfiltered CCD images taken on Nov. 4.236 (mag 16.3) and 5.083 UT (mag 16.5) with a 0.35-m reflector in the course of the U.K. Nova/Supernova Patrol. The new object is located at α = 11^{h}24^{m}40^{s}.12, δ = +38^{o}46′03′′.0 (equinox 2000.0), which is ≈ 42′′.0 west and 16′′.6 north of the center of NGC 3665. SN 2002hl is not present on Boles’ images from Apr. 4 (limiting mag 18.5), 19, or June 1, or on Palomar Sky Survey red and blue plates.

SUPERNOVA 2002hj R. Chornock, S. Jha, and A. V. Filippenko, University of California at

Berkeley; and B. Barris, University of Hawaii, report that inspection of a spectrum (range 400–1000 nm), obtained on Nov. 2.4 UT with the Keck II 10-m telescope (+ ESI), reveals that SN 2002hj (cf. IAUC 8006) is a supernova of type II. Well-developed P-Cyg lines of the H Balmer series, Fe II, O I, and Ca II are present on a blue continuum. Narrow Hα emission from the host galaxy at the position of the supernova is present at a redshift of 7090 km/s. The absorption minimum of Hβ has an expansion velocity of 8900 km/s.

RADIO TRANSIENT IN SCORPIUS O. B. Slee, G. S. Tsarevsky, and R. J. Sault, Australia Telescope Na-

tional Facility; M. P. Rupen, V. Dwahan, and A. J. Mioduszewski, National Radio Astronomy Observatory; and D. Campbell-Wilson and R. Deacon, University of Sydney, report that the microquasar candidate reported as a radio transient by Rupen et al. on IAUC 7968 was previously detected in 2000 and 2001 by the Australia Telescope Compact Array (ATCA) as a 12.3-mJy, flat-spectrum radio source and showed a high level of variability. Further ATCA, Very Large Array, and MOST-satellite observations, spanning 843 MHz to 8.6 GHz during 2002 Sept. 3–Oct. 8, show that the new flare had a flat spectrum between 4.8 and 8.6 GHz and reached its peak of 7.2 mJy near Sept. 15, dropping to below the detection level of 0.3 mJy by Oct. 8. The flux density at 843 MHz was 1.5 mJy on Sept. 17, showing that source was optically thick at low frequencies. Tsarevsky et al. (URL given on IAUC 7968) identified this source with a relatively bright star (R = 11.4) having strong, variable Hα emission. Further multi-wavelength studies of this object are requested.

2002 November 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green